OpenStudy (anonymous):

I'm stuck on this question, please help me out! Select the coordinates of two points on the line y = 8. (8, 8) and (8, 0) (8, -8) and (8, 8) (0, 8) and (8, 0) (0, 8) and (8, 8) I believe the answer is D, but I am not entirely sure.

OpenStudy (mosaic):

y = 8 is the equation of a horizontal line, parallel to the x-axis, at a distance 8 above the x-axis. No matter which point you choose on this line, the y-coordinate will always be 8. x can be anything.

OpenStudy (jdoe0001):

think about it this way y = 8 and "x" can be whatever so \(\large \begin{array}{llll} x&y \\\hline\\ -\alpha&8\\ \pm 1,000&8\\ \pm 1,000,000& 8\\ +\alpha&8 \end{array}\) 

OpenStudy (anonymous):

Thank you, I understand it now!

OpenStudy (jdoe0001):

yw